OpenStudy (anonymous):

when 2x^2/y=w+2/4 is solved for w, one equation is w=8x^2/y -2. Which of the following is an equivalent equation to find w?

OpenStudy (anonymous):

1. W=8x^2 +2y / y 2. w=8x^2 -2y / y 3. w=8x^2 -3y 4. w=8x^2 - y